# thanks: $0$-0003-0728-2533thanks: $0$-0001-5501-6008

### v1

@@ IV.4 Calibrations
Table 4: Priors and median values (with $16$-th and $84$-th percentiles) for model parameters (rows) for dark matter subhalos with different inner slopes (columns).
@@
In the semi-analytic models presented in the previous subsection, there are in total four free parameters that must be calibrated to simulations. One of these parameters is $\alpha_{\mathrm{s}}$ in the tidal stripping model, the others are the parameters $\epsilon_{\mathrm{h}}$, $f_{2}$ and $\beta_{\mathrm{h}}$ in the tidal heating model. We calibrate these parameters by comparing the predictions for $M_{\rm bound}$, $V_{\mathrm{max}}$ and $R_{\mathrm{max}}$ from our models to the results measured from simulations presented in Sec. III. To perform this calibration we define a likelihood function as:
@@
[…]
@@
Here $\widetilde{\sigma}_{M,\mathrm{i}}$ and $\widetilde{\sigma}_{V,\mathrm{i}}$ are the Poisson errors measured from simulations, and $\widetilde{\sigma}_{R,\mathrm{i}}$ is defined as half of the radial bin width used for computing $V_{\mathrm{max}}$ [^7] Note that we have performed a supersampling of the subhalo profiles, thus the bin width used here is smaller than the original radial bin width (see Sec. II.3)..
@@
For model parameters $\{\alpha_{s},\epsilon_{h},f_{2},\beta_{h}\}$ we adopt uniform priors. For $\mathcal{R}$, a loguniform prior is used. The priors and resulting median values are listed in Table 4. The posteriors of model parameters for different dark matter profiles are shown in Fig. 8. We find that for $\gamma=1.5$, the tidal stripping efficiency parameter $\alpha_{s}$ is unconstrained from above. We report the 16-th percentile in Table 4 as a conservative lower bound. In practice, this means that our model might have underestimated the tidal radius $r_{t}$ in this case such that there is not enough mass outside $r_{t}$ to be stripped. Including partial of the contribution from the centrifugal force in computing $r_{t}$, i.e. assuming a nonzero value of $\gamma_{s}$ in Eq. (34), may help improve the fitting. Nevertheless, our current model already fits the bound mass evolution very well for $\gamma=1.5$, see Fig. 10.
@@ V Conclusions and discussions
We have run high-resolution idealized simulations to study the evolution of dark matter subhalos under the tidal effects from their host. We consider a generalized dark matter halo profile controlled by three parameters $\alpha$, $\beta$ and $\gamma$ (see Eq. (8)). By changing these parameters, we can represent a dark matter profile with a flat core ($\gamma=0$) or NFW-like cuspy profile ($\gamma=1$). We have run simulations with different combinations of these parameters and found the fitting functions for the tidal track in each case. The $V_{\rm max}-R_{\rm max}$ tracks we find for different $\gamma$, the inner slope of dark matter profile, are in agreement with the previous studies by Ref. [51] expect for the case with a extremely cuspy profile, i.e. $\gamma=1.5$. We have checked the convergence of the results for $\gamma=1.5$ and found that using a global time step size leads to better converged results. Our converged results still show some differences from the one in Ref. [51]. We note that Ref. [51] uses a global, fixed time step size, similar to the global time step size in our convergence tests. We find that their time step size may be too large to obtain converged tidal track.
We have also run tests with lower subhalo concentration and a host potential that grows with time, see Appendix C. While the bound mass of subhalos in these cases evolves differently than the fiducial settings (i.e. lower concentration subhalos are more influenced by the tidal effects), the tidal tracks are only marginally affected, confirming that the tidal tracks are mostly sensitive to the bound mass fraction. Furthermore, we find that adding a Miyamoto & Nagai disk [109] and Hernquist bulge [110] potential to the host to mimic the Milky Way disk and bugle also does not have a significant impact on the tidal track—the difference from the fiducial case is less than $10\%$ (see Appendix D).
@@
For CDM, it is well known that the dark matter halos are well described by the NFW profile. But for other types of dark matter particle this is not necessarily true. For example, for self-interacting dark matter (SIDM), due to frequently scattering between dark matter particles in the halo center, a constant density core can form. At later stages of the evolution of SIDM halos, core collapse can happen leading to a very cuspy density profile. Core formation can also happen in other dark matter models such as fuzzy dark matter due to additional pressure from the quantum effects or, in CDM models, due to baryonic feedback. Thus considering the evolution for different dark matter profiles is useful to allow comparison of different dark matter models with observations of subhalos. Using the semi-analytic models presented in this work will allow us to predict the statistical properties of subhalos in different dark matter models and distinguish these models by comparing with observations such as the Milky Way satellite populations and strong gravitational lenses.
@@
Furthermore, in the current work, we have ignored non-gravitational interactions between dark matter particles. In future works, we will explore the possibility to include other effects in different dark matter scenarios, e.g. enhanced tidal stripping in fuzzy dark matter models due to “quantum tunnelling” [18, 114], core evolution [54, 55, 56, 57, 58] and ram pressure stripping [115] in SIDM model.
@@ Acknowledgements
X.D. thanks Jorge Peñarrubia for beneficial discussions on initial conditions and simulation with SUPERBOX. X.D. and T.T. acknowledge support from the National Science Foundation through grant NSF-AST-1836016 and by the Gordon and Betty Moore Foundation through grant 8548.
@@ Appendix A Evolution of $R_{\mathrm{max}}$ and $V_{\mathrm{max}}$ as functions of the bound mass fraction
Figure 13: $R_{\mathrm{max}}$ and $V_{\mathrm{max}}$ for $(\alpha,\beta)=(1,4)$ and different values of $\gamma$ from N-body simulations using GADGET-4. Solid lines show the fitting functions found in this work.
@@ Appendix C Effects of subhalo concentration and time-evolving host potential
To test how subhalo concentration may affect the tidal tracks. We run a few tests for subhalos with NFW profiles, i.e. $(\alpha,\beta,\gamma)=(1,3,1)$ and half of the fiducial concentration. As shown in Fig. 17, subhalos with lower concentrations (colored dashed curves) are more influenced by the tidal stripping and have faster mass loss compare to the fiducial cases (colored solid curve). But the $R_{\mathrm{max}}$ versus $V_{\mathrm{max}}$ tracks are only marginally affected. A detailed study of a even larger change in subhalo concentrations as done in (Green et al.) is needed to determine the possible weak dependence of tidal tracks on subhalo concentrations.
@@
Figure 17: Bound mass evolution and tidal tracks for NFW subalhos with different concentration (colored lines) and a time-evolving host potential (black line). For the case with evolving host potential, the subhalo is initially put on the same orbit as the static host case with $R_{\mathrm{p}}/R_{\mathrm{a}}=1/20$.
@@ Appendix D Effects of galactic disk and bulge
In the fiducial simulations, we assume the host have an NFW profile and neglect any possible contribution from the baryons in the halo. To test whether our results are affected by the baryonic potential, we add a Miyamoto & Nagai disk potential [109]
@@
to the static host potential. For the fiducial host mass of $10^{12}\mathrm{M}_{\odot}$, we choose the following parameters in Eqs. 51 and 52 to mimic the Milky Way disk and bulge [68, 116]:
@@
The subhalo is assumed to have an NFW profile at the beginning of the simulation. We choose the ratio of pericenter to apocenter distances to be $R_{\mathrm{p}}/R_{\mathrm{a}}=1/50$ so that the subhalo can enter within the disk radius $R_{d}$. The subhalo orbital plane is tilted by a angle of $\pi/6$ with respect to the disk plane.

### v2

@@ IV.4 Calibrations
Table 4: Priors and median values (with $16$th and $84$th percentiles) for model parameters (rows) for dark matter subhalos with different inner slopes (columns).
@@
In the semi-analytic models presented in the previous subsection, there are in total four free parameters that must be calibrated to simulations. One of these parameters is $\alpha_{\mathrm{s}}$ in the tidal stripping model, the others are the parameters $\epsilon_{\mathrm{h}}$, $f_{2}$, and $\beta_{\mathrm{h}}$ in the tidal heating model. We calibrate these parameters by comparing the predictions for $M_{\rm bound}$, $V_{\mathrm{max}}$ and $R_{\mathrm{max}}$ from our models to the results measured from simulations presented in Sec. III. To perform this calibration we define a likelihood function as
@@
where $M_{\rm bound,i}^{\mathrm{p}}$, $V_{\rm max,i}^{\mathrm{p}}$, and $R_{\rm max,i}^{\mathrm{p}}$ are model predictions at the ith snapshot. Here, $\sigma_{M,\mathrm{i}}$, $\sigma_{V,\mathrm{i}}$ and $\sigma_{R,\mathrm{i}}$ represent the combined uncertainties in the measurements and the model. Given that our models, like any models, are imperfect, we introduce a free parameter $\mathcal{R}$ that quantifies the model uncertainties and write the total uncertainties as
@@
Here $\widetilde{\sigma}_{M,\mathrm{i}}$ and $\widetilde{\sigma}_{V,\mathrm{i}}$ are the Poisson errors measured from simulations, and $\widetilde{\sigma}_{R,\mathrm{i}}$ is defined as half of the radial bin width used for computing $V_{\mathrm{max}}$ [^7] Note that we have performed a supersampling of the subhalo profiles, thus the bin width used here is smaller than the original radial bin width (see Sec. II.3).
@@
For model parameters $\{\alpha_{s},\epsilon_{h},f_{2},\beta_{h}\}$ we adopt uniform priors. For $\mathcal{R}$, a loguniform prior is used. The priors and resulting median values are listed in Table 4. The posteriors of model parameters for different dark matter profiles are shown in Fig. 8. We find that for $\gamma=1.5$, the tidal stripping efficiency parameter $\alpha_{s}$ is unconstrained from above. We report the 16th percentile in Table 4 as a conservative lower bound. In practice, this means that our model might have underestimated the tidal radius $r_{t}$ in this case such that there is not enough mass outside $r_{t}$ to be stripped. Including partial of the contribution from the centrifugal force in computing $r_{t}$, i.e. assuming a nonzero value of $\gamma_{s}$ in Eq. (34), may help improve the fitting. Nevertheless, our current model already fits the bound mass evolution very well for $\gamma=1.5$, see Fig. 10.
@@ V Conclusions and discussions
We have run high-resolution idealized simulations to study the evolution of dark matter subhalos under the tidal effects from their host. We consider a generalized dark matter halo profile controlled by three parameters $\alpha$, $\beta$, and $\gamma$ [see Eq. (8)]. By changing these parameters, we can represent a dark matter profile with a flat core ($\gamma=0$) or NFW-like cuspy profile ($\gamma=1$). We have run simulations with different combinations of these parameters and found the fitting functions for the tidal track in each case. The $V_{\rm max}-R_{\rm max}$ tracks we find for different $\gamma$, the inner slope of dark matter profile, are in agreement with the previous studies by Ref. [53] expect for the case with a extremely cuspy profile, i.e. $\gamma=1.5$. We have checked the convergence of the results for $\gamma=1.5$ and found that using a global time step size leads to better converged results. Our converged results still show some differences from the one in Ref. [53]. We note that Ref. [53] uses a global, fixed time step size, similar to the global time step size in our convergence tests. We find that their time step size may be too large to obtain converged tidal track.
We have also run tests with lower subhalo concentration and a host potential that grows with time, see Appendix C. While the bound mass of subhalos in these cases evolves differently than the fiducial settings (i.e. lower concentration subhalos are more influenced by the tidal effects), the tidal tracks are only marginally affected, confirming that the tidal tracks are mostly sensitive to the bound mass fraction. Furthermore, we find that adding a Miyamoto and Nagai disk [113] and Hernquist bulge [72] potential to the host to mimic the Milky Way disk and bugle also does not have a significant impact on the tidal track—the difference from the fiducial case is less than $10\%$ (see Appendix D).
@@
For CDM, it is well known that the dark matter halos are well described by the NFW profile. But for other types of dark matter particle this is not necessarily true. For example, for SIDM, due to frequently scattering between dark matter particles in the halo center, a constant density core can form. At later stages of the evolution of SIDM halos, core collapse can happen leading to a very cuspy density profile. Core formation can also happen in other dark matter models such as fuzzy dark matter due to additional pressure from the quantum effects or, in CDM models, due to baryonic feedback. Thus considering the evolution for different dark matter profiles is useful to allow comparison of different dark matter models with observations of subhalos. Using the semi-analytic models presented in this work will allow us to predict the statistical properties of subhalos in different dark matter models and distinguish these models by comparing with observations such as the Milky Way satellite populations and strong gravitational lenses.
@@
Furthermore, in the current work, we have ignored non-gravitational interactions between dark matter particles. In future works, we will explore the possibility to include other effects in different dark matter scenarios, e.g. enhanced tidal stripping in fuzzy dark matter models due to “quantum tunneling” [18, 117], core evolution [56, 57, 58, 59, 60] and ram pressure stripping [118] in SIDM model.
@@ Acknowledgements
X.D. thanks Jorge Peñarrubia for beneficial discussions on initial conditions and simulation with SUPERBOX. X.D. and T.T. acknowledge support from the National Science Foundation through Grant No. NSF-AST-1836016 and NSF-AST-2205100, and by the Gordon and Betty Moore Foundation through Grant No. 8548. A. N. acknowledges support from the National Science Foundation through Grant No. NSF-AST-2206315.
@@ Appendix A Evolution of $R_{\mathrm{max}}$ and $V_{\mathrm{max}}$ as functions of the bound mass fraction
Figure 13: $R_{\mathrm{max}}$ and $V_{\mathrm{max}}$ for $(\alpha,\beta)=(1,4)$ and different values of $\gamma$ from $N$-body simulations using Gadget-4. Solid black lines show the fitting functions found in this work. For comparisons, the fitting functions for the same $(\alpha,\gamma)$, but $\beta=3$ are also shown in each panel (light gray lines).
@@ Appendix C Effects of subhalo concentration and time-evolving host potential
To test how subhalo concentration may affect the tidal tracks. We run a few tests for subhalos with NFW profiles, i.e. $(\alpha,\beta,\gamma)=(1,3,1)$ and half of the fiducial concentration. As shown in Fig. 17, subhalos with lower concentrations (colored dashed curves) are more influenced by the tidal stripping and have faster mass loss compare to the fiducial cases (colored solid curve). But the $R_{\mathrm{max}}$ versus $V_{\mathrm{max}}$ tracks are only marginally affected. A detailed study of an even larger change in subhalo concentrations as done in Green et al. [54] is needed to determine the possible weak dependence of tidal tracks on subhalo concentrations.
@@
Figure 17: Bound mass evolution and tidal tracks for NFW subhalos with different concentration (colored lines) and a time-evolving host potential (black line). For the case with evolving host potential, the subhalo is initially put on the same orbit as the static host case with $R_{\mathrm{p}}/R_{\mathrm{a}}=1/20$.
@@ Appendix D Effects of galactic disk and bulge
In the fiducial simulations, we assume the host have an NFW profile and neglect any possible contribution from the baryons in the halo. To test whether our results are affected by the baryonic potential, we add a Miyamoto and Nagai disk potential [113]
@@
to the static host potential. For the fiducial host mass of $10^{12}\mathrm{M}_{\odot}$, we choose the following parameters in Eqs. (51) and (52) to mimic the Milky Way disk and bulge [71, 119]:
@@
The subhalo is assumed to have an NFW profile at the beginning of the simulation. We choose the ratio of pericenter to apocenter distances to be $R_{\mathrm{p}}/R_{\mathrm{a}}=1/50$ so that the subhalo can enter within the disk radius $R_{d}$. The subhalo orbital plane is tilted by an angle of $\pi/6$ with respect to the disk plane.